# Efficient and Universally Accessible Cross-Chain Options without Upfront Holder Collateral

### v2

@@ Abstract
In this paper, we present a protocol that addresses the aforementioned issues. Our protocol is the first to eliminate the need for holders to post collateral when establishing options in trustless service environments (i.e. without a cross-chain bridge), which is achieved by introducing a guarantee from the option writer. Its universality allows for cross-chain options involving nearly any assets on any two different blockchains, provided the chains’ programming languages can enforce and execute the necessary contract logic. Another key innovation is reducing option position transfer latency, which uses Double-Authentication-Preventing Signatures (DAPS). Our evaluation demonstrates that the proposed scheme reduces option transfer latency to less than half of that in existing methods. Rigorous security analysis proves that our protocol achieves secure option trading, even when facing adversarial behaviors.
@@ I INTRODUCTION
To trade options involving any two assets, users can use cross-chain bridges to interconnect different chains. However, current cross-chain bridges are not mature enough for widespread adoption due to numerous security risks and high expenses. According to statistics [9], cross-chain bridges have experienced at least 47 incidents, resulting in approximately 1.8 billion US dollars being hacked. Narrowly speaking, cross-chain bridges refer to protocols that enable the transfer of assets or data between different blockchain networks [10, 11, 12]. They can be categorized into “external” and “native” based on their verification methods. [^1] Some research includes optimistic verification. We categorize optimistic verification as external since it requires at least one honest watcher. External verification introduces a trusted third party (which can also be a committee) [^2] The trusted third party here can also refer to Trusted Execution Environments (TEE) for cross-chain transactions. to facilitate message transmission. This approach is vulnerable to various attacks, such as rug pulls, private key leakage [12, 9, 13, 14, 15], and software vulnerabilities [16, 17, 18]. Native verification [19, 20, 21, 22] uses light clients deployed on-chain to verify proofs. One representative solution is zkBridge [22]. However, developing complex light clients for each heterogeneous blockchain is expensive. The cross-chain verification process is also costly and complicated due to the use of zero-knowledge proofs.
Another widely used technique for cross-chain transactions is the Hashed TimeLock Contract (HTLC) [23, 24]. HTLCs do not explicitly involve forwarding assets or information from one chain to another for verification, and they only require verification by the parties involved in the transaction. Parties locally verify the state on one chain and then decide whether to proceed on the other chain. Researchers [25, 26, 27, 28, 29] have proposed using HTLCs to establish options that effectively address the limitation mentioned above: they enable the trade of any assets across any chain [^3] Here, we use “any” to stress on the wide range of chains. Indeed, it means a blockchain that has a Turing complete language for smart contracts.. HTLCs were initially designed for atomic swaps. They are shown to be equivalent to a premium-free American call option, which gives holders the right to purchase the underlying asset. [26]. However, current HTLC-based option protocols face significant challenges regarding collateral and transferability.
@@
Therefore, without holder’s upfront collateral, there is a risk that the writer could give out their assets without receiving anything in return from the holder. We can possibly utilize cross-chain bridges to solve the problem. However, as mentioned earlier, current cross-chain bridges present significant risks. Therefore, our goal is to design a protocol in a trustless service environment to facilitate option traders’ transactions.
For transferability, most research focuses on establishing the option contract between the holder and writer. Once an option is established, it should be tradable since itself is valuable. However, most works overlook this crucial feature. In [30], the authors introduced transferable cross-chain options, allowing options to be transferred post-establishment. Despite its innovation, their approach has two limitations. The first one is prolonged transfer time. In this protocol, the option transfer requires a staggered $9\Delta$ (approximately 9 hours on Bitcoin), where $\Delta$ represents the time period sufficient for each party to release, broadcast, and confirm transactions on the blockchain, typically 1 hour on Bitcoin. The second one is it is vulnerable to the Phantom Bid Attack, where an adversary creates multiple fake buyers who offer higher prices but fail to finalize the transfer, the option holder/writer is unable to sell their position.
@@
TABLE I: Comparison between our protocol and other cross-chain protocols. “Universal” represents whether or not parties are allowed to create their own assets on different chains for option trading and $\Delta$ represents the block confirmation time.
@@
Deri V4 [8] - Cross-chain application is adopted. This involves market making and taking inputs from an external oracle.
@@
Mojtaba et al. [29] - Based on HTLC, compensation is provided to the other party if it does not conform during the option establishment phase.
@@
Engel et al. [30] $9\Delta$ HTLC-based cross-chain options position transfer is proposed.
Our Protocol $4\Delta$
@@
  • Holder Collateral-Free Cross-Chain Options. We design a option protocol without requiring the holder to post collateral when the option is established. The key idea is the introduction of economic incentives. We let the option writer provide a fair guarantee before the option is established. The guarantee serves as an compensation to the holder if the writer fails to fulfill the obligation when the holder exercises the option.
  • Combing the above two schemes, we propose a protocol for efficient cross-chain options without upfront collateral from holders, achieving a option trading protocol closer to those of traditional options. What’s more, the protocol can support concurrent bidding of multiple potential buyers. The option transfer time is constant regardless of the number of bidders.
@@ II PRELIMINARIES
Blockchain, Smart Contract, and Asset. A blockchain is a tamper-proof distributed ledger that records asset balances for each address. An asset can be a cryptocurrency, a token, or any item transactable on-chain. A party can be an individual, organization, or any entity capable of interacting with the blockchain. A smart contract (simply as “contract”) is an immutable agreement written in code and executed by blockchain system. Parties can create contracts, call functions, and check contract code and state. $\Delta$ represents the sufficient time period for the parties to release, broadcast, and confirm transactions on the blockchain.
@@
In this paper, Double-Authentication-Preventing Signatures (DAPS) [31] is a key component in our protocol design. Initially, DAPS are designed to inhibit the reuse of a single private key for signing two different messages, where a message consists of a pair of message address and message payload in the form of $(a,p)$. DAPS ensures that a particular secret key $sk$ cannot sign the same address $a$ with different payloads $p$. This property can be used for preventing double spending in blockchains. Two messages $m_{1}=(a_{1},p_{1})$ and $m_{2}=(a_{2},p_{2})$ are considered colliding if $a_{1}=a_{2}$ and $p_{1}\neq p_{2}$. Any two signatures with the identical address but different contents will disclose the secret key. Given a security parameter $\lambda$, DAPS can be delineated as follows.
@@
Hashed TimeLock Contracts (HTLCs). The Hashed TimeLock Contract (HTLC) is a cryptographic contract utilized to facilitate secure and trustless transactions. In an vanilla HTLC, funds are locked in a contract and can only be accessed by the designated recipient upon fulfillment of predetermined conditions within a specified time frame $T$. The condition is expressed as the presentation of a preimage of the hash. For example, the contract asks the designated recipient to present the preimage $A$ for the hash $H(A)$. If $A$ is not provided before $T$, the funds are refunded after $T$.
@@ III OVERVIEW
Consider the scenario described in Section I. Alice and Bob strike a deal: Bob agrees to sell his 100 guilder coins to Alice in exchange for 100 florin coins within one week. To secure this arrangement, Alice pays a premium with another kind of asset, e.g. 1 US dollar, for the option to complete the transaction. In the world of blockchain, the 100 florin and 100 guilder coins are referred to as $Asset_{A}$ on $Chain_{A}$ and $Asset_{B}$ on $Chain_{B}$, respectively, with one week as the expiration time $T_{E}$. A 1 US dollar premium could be asset $P$ on another $Chain_{P}$ including $Chain_{A}$ and $Chain_{B}$. An HTLC-based option requires Alice to pay a premium $P$ and escrow $Asset_{A}$ at the beginning of the option transaction, followed by Bob escrowing $Asset_{B}$.
@@
Making transfer more robust and efficient. Consider an HTLC-based option where Alice and Bob lock their assets with the hashlock $H(A)$, where $A$ is the exercise secret generated by Alice. Now, consider the case where Carol purchases Alice’s position. In the previous work [30], the protocol first locks the contracts by tentatively assigning a new hashlock to replace the old one. Considering that Alice can place different hashlocks on two chains and Carol may replace Alice on one chain but not the other, Bob is given a time-consuming consistency phase to ensure the new hashlocks are consistent. We would like to reduce the time needed for Alice’s position transfer.
@@
We manage to reduce the transfer time by a key observation. Since Alice cannot transfer the option to multiple bidders simultaneously, it logically prompts the use of Double-Authentication-Preventing Signatures (DAPS) to prevent a seller from selling signatures to multiple bidders, thus remove the requirement of guaranteeing consistency by Bob’s efforts. If multiple signatures are revealed, a secret key can be extracted by DAPS, then punishment is enforced automatically to ensure fair payoff. By adopting DAPS, the transaction completion in our protocol is less than half the time of the previous method [30].
@@
Holder Collateral-Free Cross-Chain Options. Our objective is to allow Alice to pay a premium to secure this right, without the need to escrow the assets in advance. We need a mechanism to enable the correct exercise of this right —Alice pays her florin coins to get Bob’s guilder coins. Alice cannot get Bob’s coins without paying florin coins to Bob. A naive approach would be requiring cross-chain transaction confirmation of Alice’s escrow when Alice decides to exercise. However, every blockchain cannot directly get the cross-chain transaction confirmation information from another chain. This approach requires cross-chain bridges, which suffer from various security issues [32, 9], such as key leakage [15], smart contract vulnerabilities [14, 33], and rug pulls [34, 13]. The goal is to design a holder collateral-free option without a trusted cross-chain bridge.
@@
If we grant Alice direct access to the exercise right (or the preimage of hash, exercise secret in HTLC), then Bob’s interests is not protected, as Alice could withdraw Bob’s coins directly. To resolve this problem, we resort to economic incentives that commonly present in the DeFi markets, which are also drive forces for options. We let Bob control the exercise secret while Alice retains the right to penalize Bob. In addition to a collateral (100 guilder coins in our example) required by normal option contracts, we ask Bob to lock another valuable asset on $Chain_{A}$ as a guarantee for Alice when she wants to exercise the right. If Alice wants to exercise her right after 2 days, she locks her asset on $Chain_{A}$ (100 florin coins in our example). If Bob does not release the exercise secret, Alice will get Bob’s guarantee as compensation. Suppose this guarantee is sufficient to compensate for the expected profit of this option. This method provides Bob with incentives to cooperate when Alice wants to exercise her right.
@@
Discussion on the Guarantee. Considering the value of guarantee, i.e. $Asset_{G}$, options trading occurs due to differing asset expectations. Alice buys options for speculation or insurance, while Bob earns a premium. This option ensures that Alice gets at least $Asset_{G}$ upon exercising. For rational Alice, if her expected return and the agreed $Asset_{G}$ exceed the premium, she will purchase the option. Typically, this premium is about 2% [26] of the underlying asset. Therefore, the guarantee addition by $Asset_{G}$ is insignificant. Our work focuses more on the mechanism setup, leaving precise pricing for future work.
@@
Integration. We then integrate the efficient option transfer protocol into the holder collateral-free cross-chain options. Bob controls the exercise secret and the transfer process instead of Alice, then the processes for transferring the holder and writer are reversed. The hash of exercise secret on both chains must remain identical after Bob transfers his position to Dave. A key challenge is ensuring that honest parties incur no losses. This involves addressing potential misbehavior by any party and collusion between any two parties. To counter these, we introduce a withdrawal delay to allow Dave to retrieve assets in case Bob acts maliciously and a transfer confirmation delay to allow Alice to contest any inconsistent replacement of the hashlock.
@@ IV PROTOCOL
Due to the complexity of efficient holder collateral-free options, we elaborate on the protocol gradually. We first introduce the efficient transfer of an option in Section IV-A. Next, we outline how to achieve holder collateral-free cross-chain options in Section IV-B. Finally, we show the efficient, holder collateral-free option protocol.
@@ IV-A Efficient Option Transfer Protocol
Option Initialization. Firstly, we illustrate an efficient option transfer protocol in an HTLC-based option. Assume Alice and Bob initialize an HTLC-based option as the holder and writer respectively. In this option, Alice locks $Asset_{A}$ on $Chain_{A}$, intending to transfer it to Bob if a preimage of $H(A)$ is presented before $T_{E}+\Delta$. Bob locks $Asset_{B}$ on $Chain_{B}$, intending to transfer it to Alice if a preimage of $H(A)$ is presented before $T_{E}$, where $T_{E}$ is the expiration time of this option. Alice owns the preimage $A$. In addition, Alice performs $\text{KeyGen}(1^{\lambda})\rightarrow(pk_{A},sk_{A})$, which acts as a transfer key pair , which are used for DAPS and misbehavior detection. $pk_{A}$ is recorded in both contracts. The transfer key is used by Alice when transferring ownership to another party. A signature generated by $sk_{A}$ can be used to replace the contract holder, the hashlock, and the new transfer public key. Similarly, Bob creates a transfer key and records it on chains. Alice and Bob agree in advance on a value (e.g., a 256-bit random number) to serve as the message address $a$ recorded in the contracts for the DAPS. We take holder position transfer as an example to illustrate this transfer protocol.

### v4

# Cross-Chain Options: A Bridgeless, Universal, and Efficient Approach

@@ Abstract
In this paper, we present a protocol that addresses the aforementioned issues. Our protocol is the first to eliminate the need for holders to post collateral when establishing options in trustless service environments (i.e., without a cross-chain bridge), which is achieved by introducing a guarantee from the option writer. Its universality allows for cross-chain options involving nearly any assets on any two different blockchains, provided the chains’ programming languages can enforce and execute the necessary contract logic. Another key innovation is reducing option position transfer latency, which uses Double-Authentication-Preventing Signatures (DAPS). Our evaluation demonstrates that the proposed scheme reduces option transfer latency to less than half of that in existing methods. Rigorous security analysis proves that our protocol achieves secure option trading, even when facing adversarial behaviors.
@@ I INTRODUCTION
To trade options involving any two assets, users can use cross-chain bridges to interconnect different chains. However, current cross-chain bridges are not mature enough for widespread adoption due to numerous security risks and high expenses. According to statistics [9], cross-chain bridges have experienced at least 47 incidents, resulting in approximately 1.8 billion US dollars being hacked. Narrowly speaking, cross-chain bridges refer to protocols that enable the transfer of assets or data between different blockchain networks [10, 11, 12]. They can be categorized into “external” and “native” based on their verification methods. [^1] Some research includes optimistic verification. We categorize optimistic verification as external since it requires at least one honest watcher. External verification introduces a trusted third party [^2] The trusted third party here can also refer to Trusted Execution Environments (TEE) or a committee for cross-chain transactions. to facilitate message transmission. This approach is vulnerable to various attacks, such as rug pulls, private key leakage [12, 9, 13, 14, 15], and software vulnerabilities [16, 17, 18]. Native verification [19, 20, 21, 22] uses light clients deployed on-chain to verify proofs. One representative solution is zkBridge [22]. However, developing complex light clients for each heterogeneous blockchain is expensive. The cross-chain verification process is also costly and complicated due to the use of zero-knowledge proofs.
Another widely used technique for cross-chain transactions is the Hashed TimeLock Contract (HTLC) [23, 24]. HTLCs do not explicitly involve forwarding assets or information from one chain to another for verification, and they only require verification by the parties involved in the transaction. Parties locally verify the state on one chain and then decide whether to proceed on the other chain. Researchers [25, 26, 27, 28, 29] have proposed using HTLCs to establish options that effectively address the limitation mentioned above: they enable the trade of any assets across any chain [^3] Here, we use “any” to stress on the wide range of chains. Indeed, it means a blockchain that has a Turing-complete language for smart contracts.. HTLCs were initially designed for atomic swaps. They are shown to be equivalent to a premium-free American call option, which gives holders the right to purchase the underlying asset. [26]. However, current HTLC-based option protocols face significant challenges regarding collateral and transferability.
@@
Therefore, without the holder’s upfront collateral, there is a risk that the writer could give out their assets without receiving anything in return from the holder. We can possibly utilize cross-chain bridges to solve the problem. However, as mentioned earlier, current cross-chain bridges present significant risks. Therefore, our goal is to design a protocol in a trustless service environment to facilitate option traders’ transactions.
For transferability, most research focuses on establishing the option contract between the holder and writer. Once an option is established, it should be tradable since it is valuable. However, most works overlook this crucial feature. Previous researchers [30] introduced transferable cross-chain options (short for TCO), allowing options to be transferred post-establishment. Despite its innovation, their approach has two limitations. The first one is prolonged transfer time. In this protocol, the option transfer requires a staggered $9\Delta$ (approximately 9 hours on Bitcoin), where $\Delta$ represents the time period sufficient for each party to release, broadcast, and confirm transactions on the blockchain, typically 1 hour on Bitcoin. The second one is that it is vulnerable to the Phantom Bid Attack, where an adversary creates multiple fake buyers who offer higher prices but fail to finalize the transfer, the option holder/writer is unable to sell their position.
@@
TABLE I: Comparison between our protocol and other cross-chain protocols. “Universal” represents whether or not parties are allowed to create their own assets on different chains for option trading, and $\Delta$ represents the block confirmation time.
@@
Deri V4 [8] - Require market making and inputs from an external oracle.
@@
CFA [29] - Compensate to conforming party during option setup phase.
@@
TCO [30] $9\Delta$ HTLC-based cross-chain options position transfer is proposed.
Our Protocol $4\Delta$ Fully trustless, transferable, and efficient.
@@
  • Holder Collateral-Free Cross-Chain Options. We design an option protocol without requiring the holder to post collateral when the option is established. The key idea is the introduction of economic incentives. We let the option writer provide a fair guarantee before the option is established. The guarantee serves as a compensation to the holder if the writer fails to fulfill the obligation when the holder exercises the option.
  • Combining the above two schemes, we propose a protocol for efficient cross-chain options without upfront collateral from holders, achieving an option trading protocol closer to those of traditional options. What’s more, the protocol can support concurrent bidding of multiple potential buyers. The option transfer time is constant regardless of the number of bidders.
@@ II PRELIMINARIES
Blockchain, Smart Contract, and Asset. A blockchain is a tamper-proof distributed ledger that records asset balances for each address. An asset can be a cryptocurrency, a token, or any item transactable on-chain. A party can be an individual, organization, or any entity capable of interacting with the blockchain. A smart contract (simply as “contract”) is an immutable agreement written in code and executed by a blockchain system. Parties can create contracts, call functions, and check contract code and state. $\Delta$ represents the sufficient time period for the parties to release, broadcast, and confirm transactions on the blockchain.
@@
In this paper, Double-Authentication-Preventing Signatures (DAPS) [31] is a key component in our protocol design. Initially, DAPS is designed to inhibit the reuse of a single private key for signing two different messages, where a message consists of a pair of message address and message payload in the form of $(a,p)$. DAPS ensures that a particular secret key $sk$ cannot sign the same address $a$ with different payloads $p$. This property can be used to prevent double spending in blockchains. Two messages $m_{1}=(a_{1},p_{1})$ and $m_{2}=(a_{2},p_{2})$ are considered colliding if $a_{1}=a_{2}$ and $p_{1}\neq p_{2}$. Any two signatures with the identical address but different contents will disclose the secret key. Given a security parameter $\lambda$, DAPS can be delineated as follows.
@@
Hashed TimeLock Contracts (HTLCs). The Hashed TimeLock Contract (HTLC) is a cryptographic contract utilized to facilitate secure and trustless transactions. In a vanilla HTLC, funds are locked in a contract and can only be accessed by the designated recipient upon fulfillment of predetermined conditions within a specified time frame $T$. The condition is expressed as the presentation of a preimage of the hash. For example, the contract asks the designated recipient to present the preimage $A$ for the hash $H(A)$. If $A$ is not provided before $T$, the funds are refunded after $T$.
@@ III OVERVIEW
Consider the scenario described in Section I. Alice and Bob strike a deal: Bob agrees to sell his 100 guilder coins to Alice in exchange for 100 florin coins within one week. To secure this arrangement, Alice pays a premium with another kind of asset, e.g., 1 US dollar, for the option to complete the transaction. In the world of blockchain, the 100 florin and 100 guilder coins are referred to as $Asset_{A}$ on $Chain_{A}$ and $Asset_{B}$ on $Chain_{B}$, respectively, with one week as the expiration time $T_{E}$. A 1 US dollar premium could be an asset $P$ on another $Chain_{P}$, including $Chain_{A}$ and $Chain_{B}$. An HTLC-based option requires Alice to pay a premium $P$ and escrow $Asset_{A}$ at the beginning of the option transaction, followed by Bob escrowing $Asset_{B}$.
@@
Making transfer more robust and efficient. Consider an HTLC-based option where Alice and Bob lock their assets with the hashlock $H(A)$, where $A$ is the exercise secret generated by Alice. Consider the case where Carol purchases Alice’s position. In the previous work (TCO) [30], the protocol first locks the contracts by tentatively assigning a new hashlock to replace the old one. Alice can place different hashlocks on two chains, and Carol may replace Alice on one chain but not another. Bob is given a time-consuming consistency phase to ensure the new hashlocks are consistent. We would like to reduce the time needed for Alice’s position transfer.
@@
We manage to reduce the transfer time by a key observation. Since Alice cannot transfer the option to multiple bidders simultaneously, it logically prompts the use of Double-Authentication-Preventing Signatures (DAPS) to prevent a seller from selling signatures to multiple bidders, thus removing the requirement of guaranteeing consistency by Bob’s efforts. If multiple signatures are revealed, a secret key can be extracted by DAPS, and then punishment is enforced automatically to ensure a fair payoff. With DAPS, the transaction completion in our protocol is less than half the time of TCO [30].
@@
Holder Collateral-Free Cross-Chain Options. Our objective is to allow Alice to pay a premium to secure this right, without the need to escrow the assets in advance. We need a mechanism to enable the correct exercise of this right —Alice pays her florin coins to get Bob’s guilder coins. Alice cannot get Bob’s coins without paying florin coins to Bob. A naive approach would be to require cross-chain transaction confirmation of Alice’s escrow when Alice decides to exercise. However, every blockchain cannot directly get the cross-chain transaction confirmation information from another chain. This approach requires cross-chain bridges, which suffer from various security issues [32, 9], such as key leakage [15], smart contract vulnerabilities [14, 33], and rug pulls [34, 13]. The goal is to design a holder collateral-free option without a trusted cross-chain bridge.
@@
If we grant Alice direct access to the exercise right (exercise secret in HTLC), then Bob’s interests are not protected, as Alice could withdraw Bob’s coins directly. To resolve this problem, we resort to economic incentives that are commonly present in the DeFi markets, which are also driving forces for options. We let Bob control the exercise secret while Alice retains the right to penalize Bob. In addition to a collateral (100 guilder coins in our example) required by normal option contracts, we ask Bob to lock another valuable asset on $Chain_{A}$ as a guarantee for Alice when she wants to exercise the right. If Alice wants to exercise her right after 2 days, she locks her asset on $Chain_{A}$ (100 florin coins in our example). If Bob does not release the exercise secret, Alice will get Bob’s guarantee as compensation. Suppose this guarantee is sufficient to compensate for the expected profit of this option. This method provides Bob with incentives to cooperate when Alice wants to exercise her right.
@@
Discussion on the Guarantee. Considering the value of guarantee, i.e., $Asset_{G}$, options trading occurs due to differing asset expectations. Alice buys options for speculation or insurance, while Bob earns a premium. This option ensures that Alice gets at least $Asset_{G}$ upon exercising. For rational Alice, if her expected return and the agreed $Asset_{G}$ exceed the premium, she will purchase the option. Typically, this premium is about 2% [26] of the underlying asset. Therefore, the guarantee addition by $Asset_{G}$ is insignificant. Our work focuses on the mechanism setup, leaving precise pricing for future work.
@@
Integration. We integrate the efficient option transfer protocol into the holder collateral-free cross-chain options. Bob controls the exercise secret and transfer process instead of Alice, the processes for transferring the holder and writer are reversed. The hash of the exercise secret on the chains must remain identical after Bob transfers his position to Dave. A key challenge is ensuring that honest parties incur no loss by addressing potential misbehavior by any party and collusion between any two parties. To counter these, we introduce a withdrawal delay to allow Dave to retrieve assets in case Bob acts maliciously and a transfer confirmation delay to allow Alice to contest any inconsistent replacement of the hashlock.
@@ IV PROTOCOL
Due to the complexity of efficient holder collateral-free options, we will elaborate on the protocol gradually. We first introduce the efficient transfer of an option in Section IV-A. Next, we outline how to achieve holder collateral-free cross-chain options in Section IV-B. Finally, we show the efficient, holder collateral-free option protocol.
@@ IV-A Efficient Option Transfer Protocol
Option Initialization. Firstly, we illustrate an efficient option transfer protocol in an HTLC-based option. Assume Alice and Bob initialize an HTLC-based option as the holder and writer, respectively. In this option, Alice locks $Asset_{A}$ on $Chain_{A}$, intending to transfer it to Bob if a preimage of $H(A)$ is presented before $T_{E}+\Delta$. Bob locks $Asset_{B}$ on $Chain_{B}$, intending to transfer it to Alice if a preimage of $H(A)$ is presented before $T_{E}$, where $T_{E}$ is the expiration time of this option. Alice owns the preimage $A$.
In addition, Alice performs $\text{KeyGen}(1^{\lambda})\rightarrow(pk_{A},sk_{A})$, which acts as a transfer key pair. This transfer key pair is used as a DAPS key pair for misbehavior detection, where $pk_{A}$ is recorded in both contracts. The transfer key is used by Alice when transferring ownership to another party. A signature generated by $sk_{A}$ can be used to replace the contract holder, the hashlock, and the new transfer public key. Similarly, Bob creates a transfer key and records it on chains. Alice and Bob agree in advance on a value (e.g., a 256-bit random number) to serve as the message address $a$ recorded in the contracts for the DAPS. We take the holder position transfer as an example to illustrate this transfer protocol.